OpenStudy (anonymous):

Find (ƒ*g)(x) for the functions provided: ƒ(x) = 2*2, g(x) = (x + 1)2

OpenStudy (tanner23456):

(f*g)(x) = f(x) * g(x)

OpenStudy (solomonzelman):

you mean \(\large\color{black}{ (f\times g)(x) }\) what tanner said, or \(\large\color{black}{ (f^{~\circ } g)(x) }\) ?

OpenStudy (anonymous):

(f×g)(x) @SolomonZelman @tanner23456

OpenStudy (tanner23456):

cool so if f(x) is really 2*2 then f(x)=4 g(x) = (x+1)2 distribute the 2 and it becomes 2x+2 so f(x) * g(x) is the same thing as (4) * (2x+2) distribute the 4 and it becomes 8x + 8